### v1

@@ 8.2. Case $d\geq 2$, the upper bound
$y_{k}:=m_{X}(\varphi_{k}),\qquad y^{\prime}_{k}:=m_{X}(\varphi_{k})\log(1+m_{k}).$
@@
Replacing ˜8.8 by Proposition˜8.17 in the proof of Lemma˜8.10, we obtain the following.
@@
$Y_{k}(\boldsymbol{s})=\varphi_{k}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)},\qquad Z_{k}(\boldsymbol{s})=Y_{k}(\boldsymbol{s})-y_{k},$
Lemma˜8.11 becomes (with a similar proof) the following.
@@
We deduce the following replacement for Proposition˜8.12.
@@
$\psi(q)\geq q^{-1}\log^{-2}(q),\quad\forall q\in\operatorname{K_{small}}.$
@@
Same as for Proposition˜8.12 but using $y_{j}^{\prime}$ to define the partition $J=J_{1}\sqcup J_{2}\sqcup J_{3}$, and noting that all the upper bounds in the proof of Proposition˜8.12 are valid with $y_{j}^{\prime}$ at the place of $y_{j}$ thanks to Lemmas 8.18, 8.19, and the inequality $y_{j}^{\prime}\geq y_{j}$ (which is valid thanks to the assumption $\inf J\ggg_{\gamma_{1}}1$). ∎
We can now combine Proposition˜8.20 and Lemma˜8.13 (note here that we allow $(y^{\prime}_{j})$ to be different from $(y_{j})$ in the latter) to obtain the following.
@@
(8.33) $\lvert Z_{\operatorname{K_{small}}(n)}(\boldsymbol{s})\rvert\leq\left(y^{\prime}_{\operatorname{K_{small}}(n)}\right)^{4/5}.$
@@
(8.34) $y^{\prime}_{\operatorname{K_{small}}(n)}\leq y_{\operatorname{K_{small}}(n)}\log(1+m_{\max\operatorname{K_{small}}(n)}).$
@@
(8.35) $\forall k\in\operatorname{K_{small}},\quad\log(1+m_{k})\ll\left(y_{\operatorname{K_{small}}(k)}+O_{\gamma_{2}}(1)\right)^{1/8}.$
@@
Lemma˜8.21 and the above claim yield in particular that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for sufficiently large $n$,
@@
where $P_{k}=[0,r_{k})\times[r_{k},\tau r_{k})$. The goal is to show Lemma˜8.16 when $d=1$. Provided $\sum_{k\in\operatorname{K_{big}}}\psi(\tau^{k})\tau^{k}=+\infty$, the argument for Lemma˜8.7 yields the desired upper bound for the partial sums $\sum_{\operatorname{K_{big}}(n)}\mathscr{S}_{k}^{+}(\boldsymbol{s})$. Therefore we only need to deal with $\sum_{\operatorname{K_{small}}(n)}\mathscr{S}_{k}^{+}(\boldsymbol{s})$, and under the assumption $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})\tau^{k}=+\infty$ (as noted for Lemma˜8.16).
@@
$\varphi^{+(m)}_{k}=\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}^{(m)}_{P_{k}^{+}})\quad\text{and}\quad Y^{+(m)}_{k}(\boldsymbol{s})=\varphi^{+(m)}_{k}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
Let $(m_{k})_{k\in\operatorname{K_{small}}}$ be as in (8.31). We use the shorthand $\varphi^{+}_{k}:=\varphi^{+(m_{k})}_{k}$, $Y^{+}_{k}:=Y^{+(m_{k})}_{k}$. The argument used for the lower bound (case $d=1$) then shows that, provided $\gamma_{1}\lll 1$, we have for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$,
(8.36) $\sum_{k\in\operatorname{K_{small}}(n)}Y^{+}_{k}(\boldsymbol{s})\leq(\tau-1+10\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
We now compare the left hand side of (8.36) with $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$. In other terms, we need to show that the truncation of the cusp induced by $\chi_{k}$, and most importantly the reduction of the Siegel transform to counting $m_{k}$-primitive lattice points does not affect too much the asymptotic of the partial sums. The next lemma is a first step to replace $m_{k}$ by a term $m^{\prime}_{k}$ which grows exponentially in $k$.
@@
Let $m^{\prime}_{k}:=\max\{m_{k},t_{k}^{\gamma_{2}}\}$. Provided that $\gamma_{1},\gamma_{2}\lll 1$, we have for every $k\in\operatorname{K_{small}}$,
$\mathbb{E}\bigl{[}Y^{+(m^{\prime}_{k})}_{k}-Y^{+}_{k}\bigr{]}\ll r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}}.$
@@
We may assume $m_{k}\leq t_{k}^{\gamma_{2}}$. Unfolding definitions, then using effective single equidistribution (8.1) while noting that ${\mathcal{S}}_{\infty,l}(\varphi_{k}^{+(m^{\prime}_{k})}-\varphi_{k}^{+})\ll t_{k}^{M\gamma_{2}}$ for some $M$ as in Lemma˜8.18, we have
$\mathbb{E}\bigl{[}Y^{+(m^{\prime}_{k})}_{k}-Y^{+}_{k}\bigr{]}$ $=\int_{\mathbb{R}^{d}}\theta*[\chi_{k}(\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}^{+}}-\widetilde{{\mathbbm{1}}}^{(m_{k})}_{P_{k}^{+}})]\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}\,\mathrm{d}\xi(\boldsymbol{s})$
@@
where the last inequality uses (8.29), the definition of $c_{m_{k}}$ and assumes $\gamma_{2},\gamma_{1}$ small enough depending on $c$. ∎
We deduce that (8.36) is still valid for $Y^{+(m^{\prime}_{k})}_{k}$ in the place of $Y^{+}_{k}$.
@@
$\sum_{k\in\operatorname{K_{small}}(n)}Y^{+(m^{\prime}_{k})}_{k}(\boldsymbol{s})\leq(\tau-1+11\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
@@
Let $j\geq 1$, set $I_{j}:=\{n\in\operatorname{K_{small}}\,:\,\sum_{\operatorname{K_{small}}(n)}r_{k}^{2}\in(j,j+1]\}$. We have by Lemma˜8.22
$\mathbb{E}[\sum_{k\in I_{j}}(Y^{+(m^{\prime}_{k})}_{k}-Y^{+}_{k})]\ll\sum_{k\in I_{j}}(r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}})\leq\frac{j+1}{e^{j^{1/8}}-1}+\sum_{k\in I_{j}}t_{k}^{-\gamma_{2}},$
[…]
To conclude, we show the reduction prescribed by $m^{\prime}_{k}$ is loose enough not to affect the counting.
Lemma 8.24.
Assume $\gamma_{1}\lll\gamma_{2}$. For $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $k$, we have
$\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq Y^{+(m^{\prime}_{k})}_{k}(\boldsymbol{s}).$
Proof.
Note that we have $\theta*(\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}^{+}})\geq\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}$, because $gP^{+}_{k}\supseteq P_{k}$ for all $g\in\operatorname{supp}\theta$. Therefore, recalling the truncation function $\chi_{k}={\mathbbm{1}}_{\{\operatorname{inj}\geq t_{k}^{-\gamma_{2}}\}}$, and writing $\chi^{\prime}_{k}$ the characteristic function of the set of points $x$ such that $B_{1}x\subseteq\operatorname{supp}\chi_{k}$, we have $\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}^{+}})\geq\chi^{\prime}_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}$. It follows that
(8.37) $Y^{+(m^{\prime}_{k})}_{k}(\boldsymbol{s})\geq\bigl{(}\chi^{\prime}_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}\bigr{)}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
By Lemma˜8.5, for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $k\in\operatorname{K_{small}}$, we have $\operatorname{dist}(a(t_{k})u(\boldsymbol{s})x_{0},x_{0})\ll\log\log t_{k}$, meaning that
(8.38) $a(t_{k})u(\boldsymbol{s})\in g_{k}\operatorname{SL}_{2}(\mathbb{Z})\text{ for some $g_{k}\in G$ satisfying }\lVert g_{k}\rVert\leq(\log t_{k})^{O(1)}.$
In particular, $a(t_{k})u(\boldsymbol{s})x_{0}\in\operatorname{supp}\chi^{\prime}_{k}$ for $k$ sufficiently large, whence the truncation $\chi^{\prime}_{k}$ plays no role. Moreover, as $k\in\operatorname{K_{small}}$, we have that $P_{k}$ is included in the $t_{k}^{O(\gamma_{1})}$-neighborhood of the origin. Combined with (8.38), this yields that every point $(p,q)\in g_{k}^{-1}P_{k}\cap\mathbb{Z}^{2}$ satisfies $\operatorname{GCD}(p,q)\leq t_{k}^{\gamma_{2}}\leq m^{\prime}_{k}$ (provided $\gamma_{1}\lll\gamma_{2}$). It follows that
(8.39) $\bigl{(}\chi^{\prime}_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}\bigr{)}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}=\widetilde{{\mathbbm{1}}}_{P_{k}}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}=\mathscr{S}^{+}_{k}(\boldsymbol{s}).$
Equations (8.37) and (8.39) together finish the proof. ∎
The combination of Lemmas 8.23, 8.24 yields for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$,
@@
This concludes the proof of the asymptotic upper bound in ˜8.3 in the case $d=1$.

### v2

@@ 9.2. Case $d\geq 2$, the upper bound
$y_{k}:=m_{X}(\varphi_{k}),\hskip 23.49976pty^{\prime}_{k}:=m_{X}(\varphi_{k})\log^{2}(1+m_{k}).$
@@
Replacing ˜9.8 by Proposition˜9.17 in the proof of Lemma˜9.10, we obtain the following.
@@
$Y_{k}(\boldsymbol{s})=\varphi_{k}\bigl(a(t_{k})u(\boldsymbol{s})x_{0}\bigr),\hskip 23.49976ptZ_{k}(\boldsymbol{s})=Y_{k}(\boldsymbol{s})-y_{k},$
Lemma˜9.11 becomes (with a similar proof) the following.
@@
We deduce the following replacement for Proposition˜9.12.
@@
$\psi(q)\geq q^{-1}\log^{-2}(q),\qquad\forall q\in\operatorname{K_{small}}.$
@@
Same as for Proposition˜9.12 but using $y_{j}^{\prime}$ to define the partition $J=J_{1}\sqcup J_{2}\sqcup J_{3}$, and noting that all the upper bounds in the proof of Proposition˜9.12 are valid with $y_{j}^{\prime}$ at the place of $y_{j}$ thanks to Lemmas 9.18, 9.19, and the inequality $y_{j}^{\prime}\geq y_{j}$ (which is valid thanks to the assumption $\inf J\ggg_{\gamma_{1}}1$). ∎
We can now combine Proposition˜9.20 and Lemma˜9.13 (note here that we allow $(y^{\prime}_{j})$ to be different from $(y_{j})$ in the latter) to obtain the following.
@@
$\lvert Z_{\operatorname{K_{small}}(n)}(\boldsymbol{s})\rvert\leq\left(y^{\prime}_{\operatorname{K_{small}}(n)}\right)^{4/5}.$ (9.35)
@@
$y^{\prime}_{\operatorname{K_{small}}(n)}\leq y_{\operatorname{K_{small}}(n)}\log^{2}(1+m_{\max\operatorname{K_{small}}(n)}).$ (9.36)
@@
$\forall k\in\operatorname{K_{small}},\qquad\log^{2}(1+m_{k})\ll\left(y_{\operatorname{K_{small}}(k)}+O_{\gamma_{2}}(1)\right)^{1/8}.$ (9.37)
@@
Lemma˜9.21 and the above claim yield in particular that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for sufficiently large $n$,
@@
where $P_{k}=[0,r_{k})\times[r_{k},\tau r_{k})$. The goal is to show Lemma˜9.16 when $d=1$. Provided $\sum_{k\in\operatorname{K_{big}}}\psi(\tau^{k})\tau^{k}=+\infty$, the argument for Lemma˜9.7 yields the desired upper bound for the partial sums $\sum_{\operatorname{K_{big}}(n)}\mathscr{S}_{k}^{+}(\boldsymbol{s})$. Therefore we only need to deal with $\sum_{\operatorname{K_{small}}(n)}\mathscr{S}_{k}^{+}(\boldsymbol{s})$, and under the assumption $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})\tau^{k}=+\infty$ (as noted for Lemma˜9.16).
@@
$\varphi^{+(m)}_{k}=\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}^{(m)}_{P_{k}^{+}})\qquad\text{and}\qquad Y^{+(m)}_{k}(\boldsymbol{s})=\varphi^{+(m)}_{k}\bigl(a(t_{k})u(\boldsymbol{s})x_{0}\bigr).$
When no restriction on the $\operatorname{GCD}$ is prescribed, we simply write $\varphi^{+}_{k}:=\varphi^{+(\infty)}_{k}$, and $Y^{+}_{k}:=Y^{+(\infty)}_{k}$. Let $(m_{k})_{k\in\operatorname{K_{small}}}$ be as in (9.33). The argument used for the lower bound (case $d=1$) then shows that, provided $\gamma_{1}\lll\gamma_{2}\lll 1$, we have for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$,
$\sum_{k\in\operatorname{K_{small}}(n)}Y^{+(m_{k})}_{k}(\boldsymbol{s})\leq(\tau-1+10\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$ (9.38)
We now compare the left hand side of (9.38) with $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$. In other terms, we need to show that the truncation of the cusp induced by $\chi_{k}$, and most importantly the reduction of the Siegel transform to counting $m_{k}$-primitive lattice points does not affect too much the asymptotic of the partial sums. The next lemma is a preliminary step to suppress the $\operatorname{GCD}$ restriction due to $m_{k}$.
@@
Provided that $\gamma_{2}\lll 1$, we have for every $k\in\operatorname{K_{small}}$,
$\mathbb{E}\bigl[Y^{+}_{k}-Y^{+(m_{k})}_{k}\bigr]\ll r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}}.$
@@
Unfolding definitions, then using effective single equidistribution (9.1) while noting that ${\mathcal{S}}_{\infty,l}(\varphi_{k}^{+}-\varphi_{k}^{+(m_{k})})\ll\|\varphi_{k}^{+}\|_{L^{\infty}}\ll t_{k}^{M\gamma_{2}}$ for some $M$ as in Lemma˜9.18, we have
$\mathbb{E}\bigl[Y^{+}_{k}-Y^{+(m_{k})}_{k}\bigr$] $=\int_{\mathbb{R}^{d}}\theta*[\chi_{k}(\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}-\widetilde{{\mathbbm{1}}}^{(m_{k})}_{P_{k}^{+}})]\bigl(a(t_{k})u(\boldsymbol{s})x_{0}\bigr)\,\mathrm{d}\xi(\boldsymbol{s})$
@@
where the last inequality uses (9.31), the definition of $c_{m_{k}}$ and assumes $\gamma_{2}$ small enough depending on $c$. ∎
We deduce that (9.38) is still valid for $Y^{+}_{k}$ in the place of $Y^{+(m_{k})}_{k}$.
@@
$\sum_{k\in\operatorname{K_{small}}(n)}Y^{+}_{k}(\boldsymbol{s})\leq(\tau-1+11\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
@@
Let $n\geq 1$, set $I_{n}:=\{k\in\operatorname{K_{small}}\,:\,\sum_{\operatorname{K_{small}}(k)}r_{j}^{2}\in(n,n+1]\}$. We have by Lemma˜9.22
$\mathbb{E}\Bigl[\sum_{k\in I_{n}}\bigl(Y^{+}_{k}-Y^{+(m_{k})}_{k}\bigr)\Bigr]\ll\sum_{k\in I_{n}}\bigl(r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}}\bigr)\leq\frac{n+1}{e^{n^{1/16}}-1}+\sum_{k\in I_{n}}t_{k}^{-\gamma_{2}},$
where the last inequality follows from the definitions of $m_{k}$ in (9.33) and $I_{n}$. Therefore, the right hand side is summable over $n\geq 1$. Hence, for $\xi$-almost every $\boldsymbol{s}$, the total sum $\sum_{k\in\operatorname{K_{small}}}(Y^{+}_{k}-Y^{+(m_{k})}_{k})$ is $\xi$-almost-surely finite. Then the result follows from (9.38). ∎
It remains to check the restriction induced by the truncation $\chi_{k}$ can also be ignored. This can be done exactly as for the $d\geq 2$ case (by the mean of Lemma˜9.5 controlling excursions in the cusp). We get that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $k$, we have
$\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq Y^{+}_{k}(\boldsymbol{s}).$ (9.39)
The combination of Lemma˜9.23 and Equation (9.39) yields for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$,
@@
This shows Lemma˜9.16 holds for $d=1$, and concludes the proof of the asymptotic upper bound in ˜9.3 in the $1$-dimensional case.